# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 2.2. On the hull and the correcting curve
Denote $\psi(t):=\beta(t)\varphi(t),t\geq 0$, where $\beta$ is a correcting curve of $\varphi$. Let $L\leq\tilde{H}_{\varphi}$ be the Zariski closure of the subgroup generated by all one-parameter unipotent subgroups of $\tilde{H}_{\varphi}$. We will now show that the image of $\varphi$ in $\tilde{H}_{\varphi}/L$ is bounded, and since $L$ is observable [Observable_group_grosshans, Corollary 2.8], we will get that $\tilde{H}_{\varphi}=L$, which proves the claim.
Now, note that $L$ is a normal subgroup, so that $\tilde{H}_{\varphi}/L$ is an affine algebraic group. Since $L$ contains the unipotent radical of $\tilde{H}_{\varphi}$, we get that $\tilde{H}_{\varphi}/L$ is reductive. Then, there is an almost direct product decomposition $\tilde{H}_{\varphi}/L=ZD$, where $Z$ is the center, which is a torus, and $D$ is the derived group, which is a normal semi-simple group. Notice that each simple factor of $D$ must be compact since a non-compact simple group is generated by unipotents. Namely, $D$ is compact. Let $q:\tilde{H}_{\varphi}\to\tilde{H}_{\varphi}/L$ be the natural quotient map. To finish the proof, we show below that for every character $\chi:\tilde{H}_{\varphi}/L\to\mathbb{R}^{\times}$ it holds that $\chi\circ q(\psi(t))=1,\forall t\geq 0$. This will complete the proof since then $q\circ\psi([0,\infty))\subseteq Z_{a}D$, where $Z_{a}\leq Z$ is the maximal anisotropic compact torus subgroup of $Z$. Let $\chi:\tilde{H}_{\varphi}/L\to\mathbb{R}^{\times}$ be a character, and consider the representation of $\tilde{H}_{\varphi}/L$ on $\mathbb{R}^{2}$ defined by
$\overline{h}.(x,y):=\left(\chi(\overline{h})x,\chi(\overline{h})^{-1}y\right),~{}(x,y)\in\mathbb{R}^{2},~{}\overline{h}\in\tilde{H}_{\varphi}/L.$
Then $\tilde{H}_{\varphi}$ acts on $\mathbb{R}^{2}$ be pre-composing with $q$. Namely, $h.(x,y):=q(h).(x,y)$. Since $\tilde{H}_{\varphi}\leq G$ is observable, the above representation comes from a representation of $G$, see [Observable_groups_over_fields, Theorem 9]. Now, fix two non-zero real numbers $x_{0},y_{0}$. Because $\psi$ is non-contracting, we get that $\psi(t).(x_{0},0)$ and $\psi(t).(0,y_{0})$ are bounded away from 00. But this is only possible if $\psi(t).(x_{0},y_{0})$ is bounded. So by Corollary 2.13 we obtain that $\psi(t).(x_{0},y_{0})=(x_{0},y_{0})$. Thus, $\chi\circ q(\psi(t))=1,~{}\forall t\geq 0$. ∎
2.4. The non-contraction property under homomorphisms
In this section we show that the homomorphic image of a definable non-contracting curve is non-contracting, Lemma 2.15, and also we show that if a definable curve in a homomorphic image is non-contracting , then it has a definable non-contracting lift, Lemma 2.19. Finally, we establish that the non-contracting property is intrinsic for curves in groups whose radical is unipotent, Proposition 2.20.
@@ 2.4. The non-contraction property under homomorphisms
Let $G_{1},G_{2}$ be the real points of two affine algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a homomorphism of algebraic groups. Let $\varphi:[0,\infty)\to G_{1}$ be a continuous, non-contracting curve in $G_{1}$ definable in $\mathscr{S}$. Then $f\circ\varphi$ is a continuous, non-contracting curve in $G_{2}$ definable in $\mathscr{S}$.
@@
Let $G:=\mathbf{G}(\mathbb{R})$ and suppose that $U\leq G$ is a normal, unipotent algebraic subgroup. Let $q:G\to G/U$ be the natural map. Then, a curve $\varphi:[0,\infty)\to G$ is not contracting in $G$ if and only if $q\circ\varphi$ is not contracting in $G/U$.
@@
In view of Lemma 2.15, it suffices to show that if $q\circ\varphi$ is non-contracting, then $\varphi$ is non-contracting. Suppose for contradiction that $q\circ\varphi$ is non-contracting, but $\varphi$ is contracting. Let $V$ be a rational representation of the smallest dimension such that there exists a $v\in V\smallsetminus\{0\}$ with $\lim_{t\to\infty}\varphi(t).v=0$. Since $U$ is a unipotent group, the subspace of $U$-fixed points $V^{U}:=\{x\in V:u.x=x,\forall u\in U\}$ is of positive dimension. Since $U$ is a normal subgroup, $V^{U}$ is $G$-invariant. Then $G$ acts on $V/V^{U}$, and the natural quotient map $\pi:V\to V/V^{U}$ is $G$-equivariant. Now,
$\lim_{t\to\infty}\varphi(t).\pi(v)=\lim_{t\to\infty}\pi(\varphi(t).v)=0.$
Since $\dim(V/V^{U})<\dim(V)$, we get that $\pi(v)=0$. Namely, $v\in V^{U}$, and again, since $V$ is of smallest dimension with a $\varphi(t)$-contracting nonzero vector, we get $V=V^{U}$. Since $U$ acts trivially on $V$, the action of $G$ on $V$ factors through the action of $G/U$. Therefore, $\varphi(t)v=(q\circ\varphi(t))v$. Since $\varphi(t)v\to 0$ as $t\to\infty$, and $q\circ\varphi$ is non-contracting, we get a contradiction. ∎
@@
Let $G_{1},G_{2}$ be the real points of algebraic groups over $\mathbb{R}$. Suppose that $f:G_{1}\to G_{2}$ is a homomorphism and $\varphi:[0,\infty)\to f(G_{1})$ is a curve definable in an o-minimal structure $\mathscr{S}$. Then there exists a definable lift $\psi:[0,\infty)\to G_{1}$. Namely, $\psi$ is definable in $\mathscr{S}$, and $f\circ\psi=\varphi$.
@@
By the choice function theorem [Geom_cat_and_o_min_str, Theorem 4.5], there exists a definable curve $\psi:[0,\infty)\to G_{1}$ such that $(t,\psi(t))\in\mathcal{D}$ for all $t\geq 0$. ∎
@@
Let $G_{1}$, $G_{2}$ be the $\mathbb{R}$-points of irreducible semi-simple groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be homomorphism with finite kernel. Let $\psi:[0,\infty)\to G_{1}$ be a definable curve. Suppose that $f\circ\psi$ is non-contracting for $G_{2}$, then $\psi$ is non-contracting for $G_{1}$.
@@
Let $V$ be a finite dimensional irreducible rational representation of $G_{1}$. Then, by Schur’s lemma the center of $G_{1}$ acts via a character on $V\otimes\mathbb{C}$. Then, $\ker f$, which is central in $G_{1}$, acts trivially on the $n$-fold tensor product $W=\otimes^{n}(V\otimes\mathbb{C})$, where $n$ is the order of $\ker f$. Therefore, the action of $G_{1}$ on $W$ factors through an action of $G_{1}/(\ker f)=G_{2}$ on $W$. Let $v\in V$. If $\psi(t).v\to 0$, then $f\circ\psi(t).(v\otimes\cdots\otimes v)\to 0$ as $t\to\infty$. But, since $f\circ\psi$ is non-contracting, we get that $v=0$. ∎
@@
Let $G_{1},G_{2}$ be the real points of two affine algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a surjective homomorphism of algebraic groups. Let $\varphi:[0,\infty)\to G_{2}$ be a non-contracting curve in $G_{2}$ definable in $\mathscr{S}$. Then, $\varphi$ has a definable lift $\psi:[0,\infty)\to G_{1}$ that is a non-contracting in $G_{1}$.
@@
Without loss of generality, since the hull is generated by one-parameter unipotent subgroups, by modifying $\varphi$ with a bounded correcting curve and replacing $G_{2}$ by $(G_{2})_{u}$, the subgroup of $G_{2}$ generated by unipotent one-parameter subgroups, we may assume that the radical of $G_{2}$ is unipotent, call it $U_{2}$. Let $\bar{G}_{2}=G_{2}/U_{2}$ and $q_{2}:G_{2}\to\bar{G}_{2}$ be the quotient homomorphism. Let $U_{1}$ be the unipotent radical of $G_{1}$. Then, there exists a reductive subgroup $M_{1}$ of $G_{1}$ such that $G_{1}=M_{1}U_{1}$ and $M_{1}\cap U_{1}=\{e\}$, see [RA72, Page 11]. Since $f$ is surjective, $f(U_{1})$ is a unipotent normal subgroup of $G_{2}$, and hence $f(U_{1})\subset U_{2}$. In fact, $f(U_{1})=U_{2}$. This can be proved as follows. $f^{-1}(U_{2})=(M_{1}\cap f^{-1}(U_{2}))U_{1}$. Since $f^{-1}(U_{2})$ is a normal subgroup of $G_{1}$, $M_{1}\cap f^{-1}(U_{2})$ is reductive. So $f(M_{1}\cap f^{-1}(U_{2}))$ being a reductive subgroup of $U_{2}$, it is trivial. Since $f$ is surjective, we get $f(U_{1})=U_{2}$. Therefore, $q_{2}\circ f:M_{1}\to\bar{G}_{2}$ is surjective. Since $\bar{G}_{2}$ is semisimple, $q_{2}\circ f([M_{1},M_{1}])=\bar{G}_{2}$. Let $L_{2}=[M_{1},M_{1}]\cap\ker(q_{2}\circ f)$. Since $L_{2}$ is a normal subgroup of the semisimple group $[M_{1},M_{1}]$, there exists a semisimple normal subgroup $L_{1}$ of $[M_{1},M_{1}]$ such that $[M_{1},M_{1}]=L_{1}L_{2}$, where $L_{1}\cap L_{2}$ is a finite central subgroup of $L_{1}$. Therefore, $q_{2}\circ f:L_{1}\to\bar{G}_{2}$ is a surjective homomorphism whose kernel is $L_{1}\cap L_{2}$. Also, since $\bar{G}_{2}$ is generated by unipotent one-parameter subgroups, it has no compact simple normal subgroup. Therefore, $L_{1}$ also does not have a compact normal subgroup, and $L_{1}$ is generated by unipotent one-parameter subgroups.
@@
Let $\tilde{G}_{1}=L_{1}U_{1}$. Since $q_{2}\circ f(L_{1})=G_{2}/U_{2}$ and $f(U_{1})=U_{2}$, we have $f:\tilde{G}_{1}\to G_{2}$ is a surjective map. Therefore, by Lemma 2.17, there exists a definable $\psi:[0,\infty)\to\tilde{G}_{1}$ such that $f\circ\psi=\varphi$. Let $q_{1}:\tilde{G}_{1}\to\tilde{G}_{1}/U_{1}$ be the natural quotient map. Since $q_{1}:L_{1}\to\tilde{G}_{1}/U_{1}$ is surjective (it is an isomorphism), let $\tilde{\psi}:[0,\infty)\to L_{1}$ be a definable curve such that $q_{1}\circ\tilde{\psi}=q_{1}\circ\psi$. Since, $U_{1}\subset\ker q_{2}\circ f$, $q_{2}\circ f$ factors through $q_{1}$. Therefore, we have $(q_{2}\circ f)\circ\tilde{\psi}=(q_{2}\circ f)\circ\psi=q_{2}\circ\varphi$. Now $q_{2}\circ f:L_{1}\to\bar{G}_{2}$ is surjective with finite central kernel, and $q_{2}\circ\varphi$ is definable and non-contracting in $\bar{G}_{2}$. Therefore, by Lemma 2.18, $\tilde{\psi}$ is non-contracting in $L_{1}$. Therefore, $q_{1}\circ\psi=q_{1}\circ\tilde{\psi}$ is non-contracting in $\tilde{G}_{1}/U_{1}$. Therefore, by Lemma 2.16, $\psi$ is non-contracting in $\tilde{G}_{1}$. Finally, by Lemma 2.15 we get that $\psi$ is non-contracting in $G_{1}$. ∎
The following observation extends Lemma 2.18. Notice that Proposition 1.7 from the introduction is a direct consequence of the proposition below.
@@
Now suppose that $f\circ\varphi$ is non-contracting in $G_{2}$. Since $G_{1}$ is generated by unipotent one-parameter subgroups, $f(G_{1})$ is a subgroup of $G_{2}$ generated by unipotent one-parameter subgroups. Therefore, the solvable radical of $f(G_{1})$ is unipotent. Hence, $f(G_{1})$ is an observable subgroup of $G_{2}$, see [Observable_group_grosshans, Corollary 2.8]. Therefore, given any finite-dimensional rational representation of $V$ of $f(G_{1})$, $V$ is a subrepresentation of a finite-dimensional representation $W$ of $G_{2}$ restricted to $f(G_{1})$. Let $v\in V$. Suppose that $f\circ\varphi(t)v\to 0$. Since $v\in V\subset W$ and $f\circ\varphi$ is non-contracting in $G_{2}$, we conclude that $v=0$. This proves that $f\circ\varphi$ is non-contracting for $f(G_{1})$. Therefore, replacing $G_{2}$ by $f(G_{1})$, we may assume that $f$ is surjective. Therefore, by Lemma 2.19 there exists a lift $\psi:[0,\infty)\to G_{1}$ of $f\circ\varphi$ which is definable and non-contracting in $G_{1}$. By the Monotonicty Theorem [Geom_cat_and_o_min_str, Theorem 4.1], there is a $T_{0}\geq 0$ such that $\varphi$ and $\psi$ are continuous on $[T_{0},\infty)$. Let $z=\varphi(T_{0})\psi(T_{0})^{-1}\in\ker f$. So, $t\mapsto z\psi(t)$ and $t\mapsto\varphi(t)$ are two lifts of $f\circ\varphi$ from $[T_{0},\infty)$ to $G_{1}$ from the point $\varphi(T_{0})\in G_{1}$. Since $\ker f$ is finite, $f$ is a covering map, we get $\varphi(t)=z\psi(t)$ for all $t\in[0,\infty)$. Since $\psi$ is non-contracting, so is $\varphi$. ∎

### v2

@@ 2.2. On the hull and the correcting curve
Let $\psi(t):=\beta(t)\varphi(t)$ for all $t\geq 0$, where $\beta$ is a correcting curve of $\varphi$. Let $L\subseteq\tilde{H}_{\varphi}$ be the Zariski closure of the subgroup generated by all one-parameter unipotent subgroups of $\tilde{H}_{\varphi}$. Then, $L$ is observable in $G$, see Remark 2.4. We will now show that the image of $\varphi$ in $\tilde{H}_{\varphi}/L$ is bounded, which implies that $\tilde{H}_{\varphi}=L$, and proves the claim, see Remark 2.7.
Since $L$ is a normal subgroup of $\tilde{H}_{\varphi}$, we get that $\tilde{H}_{\varphi}/L$ is an affine algebraic group. Moreover, since $L$ contains the unipotent radical of $\tilde{H}_{\varphi}$, we get that $\tilde{H}_{\varphi}/L$ is reductive. Then, there is an almost direct product decomposition $\tilde{H}_{\varphi}/L=ZD$, where $Z$ is the center, which is a torus, and $D$ is the derived group, which is a normal semi-simple subgroup. Notice that each simple factor of $D$ must be compact since a non-compact simple group is generated by unipotent one-parameter subgroups. Namely, $D$ is compact. Let $q:\tilde{H}_{\varphi}\to\tilde{H}_{\varphi}/L$ be the natural quotient map. To finish the proof, we show below that for every character $\chi:\tilde{H}_{\varphi}/L\to\mathbb{R}^{\times}$ it holds that $\chi\circ q(\psi(t))=1$ for all $t\geq 0$. This will complete the proof since then $q\circ\psi([0,\infty))\subseteq Z_{a}D$, where $Z_{a}\subseteq Z$ is the maximal anisotropic compact torus subgroup of $Z$.
Let $\chi:\tilde{H}_{\varphi}/L\to\mathbb{R}^{\times}$ be a character, and consider the representation of $\tilde{H}_{\varphi}$ on $\mathbb{R}^{2}$ defined by
(2.4) $h\cdot(x,y):=\left(\chi(q(h))x,\chi(q(h))^{-1}y\right),\,\forall(x,y)\in\mathbb{R}^{2},\,\forall h\in\tilde{H}_{\varphi}.$
Since $\tilde{H}_{\varphi}\subseteq G$ is observable, the above representation extends to a representation of $G$, see [Observable_groups_over_fields, Theorem 9]. Now, pick $x_{0},y_{0}\in\mathbb{R}\setminus\{0\}$. Because $\psi$ is non-contracting, we get that $\psi(t)\cdot(x_{0},0)$ and $\psi(t)\cdot(0,y_{0})$ are bounded away from 00. In view of (2.4), we deduce that $\psi(t)\cdot(x_{0},y_{0})$ is bounded. So, by Corollary 2.15, $\psi(t)\cdot(x_{0},y_{0})=(x_{0},y_{0})$. Thus, $\chi\circ q(\psi(t))=1$ for all $t\geq 0$. ∎
2.4. The non-contracting property under homomorphisms
In this subsection, we show that the homomorphic image of a definable non/contracting curve is non/contracting, Lemma 2.17, and also we show that if a definable curve in a homomorphic image is non/contracting, then it has a definable non/contracting lift, Lemma 2.21. Finally, we establish that the non/contracting property is intrinsic for curves in groups whose radical is unipotent, Proposition 2.22.
@@ 2.4. The non-contracting property under homomorphisms
Let $G_{1}$ and $G_{2}$ be the real points of affine algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a homomorphism of algebraic groups. Let $\varphi:[0,\infty)\to G_{1}$ be a continuous, non-contracting curve in $G_{1}$ definable in $\mathscr{S}$. Then $f\circ\varphi$ is a continuous, non-contracting curve in $G_{2}$ definable in $\mathscr{S}$.
@@
Suppose that $U\subseteq G$ is a normal, unipotent algebraic subgroup. Let $q:G\to G/U$ be the natural map. Then, a $\mathscr{S}$-definable curve $\varphi:[0,\infty)\to G$ is non-contracting in $G$ if and only if $q\circ\varphi$ is non-contracting in $G/U$.
@@
By Lemma 2.17, it suffices to prove that if $q\circ\varphi$ is non-contracting, then $\varphi$ is non-contracting. Suppose for contradiction that $q\circ\varphi$ is non-contracting and $\varphi$ is contracting. Let $V$ be a rational representation of $G$ of the smallest dimension such that there exists a $v\in V\setminus\{0\}$ with $\lim_{t\to\infty}\varphi(t)\cdot v=0$. Since $U$ is a unipotent group, the subspace of $U$-fixed points $V^{U}:=\{x\in V:u\cdot x=x,\forall u\in U\}$ is of positive dimension. Since $U$ is a normal subgroup, $V^{U}$ is $G$-invariant. Then $G$ acts on $V/V^{U}$, and the natural quotient map $\pi:V\to V/V^{U}$ is $G$-equivariant. Now,
$\lim_{t\to\infty}\varphi(t)\cdot\pi(v)=\lim_{t\to\infty}\pi(\varphi(t)\cdot v)=0.$
Since $\dim(V/V^{U})<\dim(V)$, we get that $\pi(v)=0$. Namely, $v\in V^{U}$, and again, since $V$ is of smallest dimension with a $\varphi(t)$-contracting nonzero vector, we get $V=V^{U}$. Since $U$ acts trivially on $V$, the action of $G$ on $V$ factors through the action of $G/U$. Therefore, $\varphi(t)\cdot v=(q\circ\varphi(t))\cdot v$. Since $\varphi(t)\cdot v\to 0$ as $t\to\infty$, and $q\circ\varphi$ is non-contracting, we get a contradiction. ∎
@@
Let $G_{1}$ and $G_{2}$ be the real points of algebraic groups over $\mathbb{R}$. Suppose that $f:G_{1}\to G_{2}$ is a homomorphism of algebraic groups, and $\varphi:[0,\infty)\to f(G_{1})$ is a curve definable in $\mathscr{S}$. Then there exists a definable lift $\psi:[0,\infty)\to G_{1}$ of $\varphi$. Namely, $\psi$ is definable in $\mathscr{S}$, and $f\circ\psi=\varphi$.
@@
By the choice function theorem (see Theorem B.7), there exists a definable curve $\psi:[0,\infty)\to G_{1}$ such that $(t,\psi(t))\in\mathcal{D}$ for all $t\geq 0$. ∎
@@
Let $G_{1}$ and $G_{2}$ be the $\mathbb{R}$-points of semi-simple connected algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a surjective homomorphism with finite kernel. Let $\psi:[0,\infty)\to G_{1}$ be a definable curve. Suppose that $f\circ\psi$ is non-contracting for $G_{2}$, then $\psi$ is non-contracting for $G_{1}$.
@@
Let $V$ be a finite-dimensional irreducible rational representation of $G_{1}$. By Schur’s lemma, the center of $G_{1}$ acts via a character on $V\otimes\mathbb{C}$. Then, $\ker f$, which is central in $G_{1}$, acts trivially on the $n$-fold tensor product $W=\otimes^{n}(V\otimes\mathbb{C})$, where $n$ is the order of $\ker f$. Therefore, the action of $G_{1}$ on $W$ factors through an action of $G_{1}/(\ker f)\cong G_{2}$ on $W$. Let $v\in V$. If $\psi(t)\cdot v\to 0$, then $f\circ\psi(t)\cdot(v\otimes\cdots\otimes v)\to 0$ as $t\to\infty$. But, since $f\circ\psi$ is non-contracting, we get that $v=0$. ∎
@@
Let $G_{1}$ and $G_{2}$ be the real points of affine algebraic groups over $\mathbb{R}$, and let $f:G_{1}\to G_{2}$ be a surjective homomorphism of algebraic groups. Let $\varphi:[0,\infty)\to G_{2}$ be a non-contracting curve in $G_{2}$ definable in $\mathscr{S}$. Then, $\varphi$ has a $\mathscr{S}$-definable lift $\psi:[0,\infty)\to G_{1}$ which is non-contracting in $G_{1}$.
@@
Since the hull is generated by one-parameter unipotent subgroups, by modifying $\varphi$ with a bounded correcting curve and replacing $G_{2}$ by $(G_{2})_{u}$, we may assume without loss of generality that the radical of $G_{2}$ is unipotent, and we denote it by $U_{2}$. Let $\bar{G}_{2}=G_{2}/U_{2}$ and $q_{2}:G_{2}\to\bar{G}_{2}$ be the quotient homomorphism. Let $U_{1}$ be the unipotent radical of $G_{1}$. Then, there exists a reductive subgroup $M_{1}$ of $G_{1}$ such that $G_{1}=M_{1}U_{1}$ and $M_{1}\cap U_{1}=\{e\}$, see [RA72, Page 11]. Since $f$ is surjective, $f(U_{1})$ is a unipotent normal subgroup of $G_{2}$, and hence $f(U_{1})\subseteq U_{2}$. In fact, $f(U_{1})=U_{2}$. This can be proved as follows. $f^{-1}(U_{2})=(M_{1}\cap f^{-1}(U_{2}))U_{1}$. Since $f^{-1}(U_{2})$ is a normal subgroup of $G_{1}$, $M_{1}\cap f^{-1}(U_{2})$ is reductive. So $f(M_{1}\cap f^{-1}(U_{2}))$ being a reductive subgroup of $U_{2}$, it is trivial. Since $f$ is surjective, we get $f(U_{1})=U_{2}$. Therefore, $q_{2}\circ f:M_{1}\to\bar{G}_{2}$ is surjective. Since $\bar{G}_{2}$ is semisimple, $q_{2}\circ f([M_{1},M_{1}])=\bar{G}_{2}$. Let $L_{2}=[M_{1},M_{1}]\cap\ker(q_{2}\circ f)$. Since $L_{2}$ is a normal subgroup of the semisimple group $[M_{1},M_{1}]$, there exists a semisimple normal subgroup $L_{1}$ of $[M_{1},M_{1}]$ such that $[M_{1},M_{1}]=L_{1}L_{2}$, where $L_{1}\cap L_{2}$ is a finite central subgroup of $L_{1}$. Therefore, $q_{2}\circ f:L_{1}\to\bar{G}_{2}$ is a surjective homomorphism whose kernel is $L_{1}\cap L_{2}$. Also, since $\bar{G}_{2}$ is generated by unipotent one-parameter subgroups, it has no compact simple normal subgroup. Therefore, $L_{1}$ also does not have a compact normal subgroup, and $L_{1}$ is generated by unipotent one-parameter subgroups.
@@
Let $\tilde{G}_{1}=L_{1}U_{1}$. Since $q_{2}\circ f(L_{1})=G_{2}/U_{2}$ and $f(U_{1})=U_{2}$, we have $f:\tilde{G}_{1}\to G_{2}$ is a surjective map. Therefore, by Lemma 2.19, there exists a definable $\psi:[0,\infty)\to\tilde{G}_{1}$ such that $f\circ\psi=\varphi$. Let $q_{1}:\tilde{G}_{1}\to\tilde{G}_{1}/U_{1}$ be the natural quotient map. Since $q_{1}:L_{1}\to\tilde{G}_{1}/U_{1}$ is surjective (it is an isomorphism), let $\tilde{\psi}:[0,\infty)\to L_{1}$ be a definable curve such that $q_{1}\circ\tilde{\psi}=q_{1}\circ\psi$. Since $U_{1}\subseteq\ker(q_{2}\circ f)$, $q_{2}\circ f$ factors through $q_{1}$. Therefore, we have $(q_{2}\circ f)\circ\tilde{\psi}=(q_{2}\circ f)\circ\psi=q_{2}\circ\varphi$. Now $q_{2}\circ f:L_{1}\to\bar{G}_{2}$ is surjective with finite central kernel, and $q_{2}\circ\varphi$ is definable and non-contracting in $\bar{G}_{2}$. Therefore, by Lemma 2.20, $\tilde{\psi}$ is non-contracting in $L_{1}$. Therefore, $q_{1}\circ\psi=q_{1}\circ\tilde{\psi}$ is non-contracting in $\tilde{G}_{1}/U_{1}$. Therefore, by Lemma 2.18, $\psi$ is non-contracting in $\tilde{G}_{1}$. Finally, by Lemma 2.17 we get that $\psi$ is non-contracting in $G_{1}$. ∎
The following observation extends Lemma 2.20.
@@
Now suppose that $f\circ\varphi$ is non-contracting in $G_{2}$. Since $G_{1}$ is generated by unipotent one-parameter subgroups, $f(G_{1})$ is a subgroup of $G_{2}$ generated by unipotent one-parameter subgroups. Therefore, the solvable radical of $f(G_{1})$ is unipotent. Hence, $f(G_{1})$ is an observable subgroup of $G_{2}$, see [Observable_group_grosshans, Corollary 2.8]. Therefore, given any finite-dimensional rational representation of $V$ of $f(G_{1})$, $V$ is a subrepresentation of a finite-dimensional representation $W$ of $G_{2}$ restricted to $f(G_{1})$. Let $v\in V$. Suppose that $f\circ\varphi(t)\cdot v\to 0$. Since $v\in V\subseteq W$ and $f\circ\varphi$ is non-contracting in $G_{2}$, we conclude that $v=0$. This proves that $f\circ\varphi$ is non-contracting for $f(G_{1})$. Therefore, replacing $G_{2}$ by $f(G_{1})$, we may assume that $f$ is surjective. Therefore, by Lemma 2.21 there exists a lift $\psi:[0,\infty)\to G_{1}$ of $f\circ\varphi$ which is definable and non-contracting in $G_{1}$. By the monotonicty theorem (see Theorem B.5), there is a $T_{0}\geq 0$ such that $\varphi$ and $\psi$ are continuous on $[T_{0},\infty)$. Let $z=\varphi(T_{0})\psi(T_{0})^{-1}\in\ker f$. So, $t\mapsto z\psi(t)$ and $t\mapsto\varphi(t)$ are two lifts of $f\circ\varphi$ from $[T_{0},\infty)$ to $G_{1}$ from the point $\varphi(T_{0})\in G_{1}$. Since $\ker f$ is finite, $f$ is a covering map, we get $\varphi(t)=z\psi(t)$ for all $t\in[0,\infty)$. Since $\psi$ is non-contracting, so is $\varphi$. ∎